OpenStudy (anonymous):

Metal: Aluminum Mass of metal: 27.776 g Volume of water in the calorimeter: 26.0 mL Initial temperature of water in calorimeter: 25.3 °C Temperature of hot water and metal in hot water bath: 100.5 °C Final temperature reached in the calorimeter: 31.6 °C 1. Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. The water has absorbed the heat of the metal. So, qwater = qmetal q = m x c x delta t

OpenStudy (aaronq):

How much cool water and how much hot water did you have?

OpenStudy (aaronq):

Actually, it doesn't matter because you're only calculating the change in heat of the surroundings. just use the formula: q=m*C*(Tf-Ti)

OpenStudy (anonymous):

i dont think it matters

OpenStudy (anonymous):

i just need help setting up the equation

OpenStudy (aaronq):

(m) Volume of water in the calorimeter: 26.0 mL -> 1mL=1g (b/c density is 1g/mL) (Ti) Initial temperature of water in calorimeter: 25.3 °C (Tf) Final temperature reached in the calorimeter: 31.6 °C (C) specific heat capacity of water is 4.18 J / (g × °C)

OpenStudy (aaronq):

thats all the info you need, the rest (like the mass of the metal) is to confuse you

OpenStudy (anonymous):

ok thank you so much this really helps

OpenStudy (aaronq):

no problem

OpenStudy (anonymous):

Using the formula qmetal = m × c × ΔT, calculate the specific heat of the metal. ^ what about this?

OpenStudy (aaronq):

you know that q(water=)-q(metal) because the assumption is made that no heat was lost to other surroundings (like the room, your hands, etc) use the q you found in the previous part and the change in temperature of the metal and it's mass. q=m*C*(Tf-Ti)

OpenStudy (anonymous):

so that whole first part i got 685, i just dont know what to do with the whole q(water=)-q(metal)

OpenStudy (aaronq):

q(metal) is positive because heat was lost, q(water) is negative because heat was gained. so for q(water)=685 , then q(metal)=-685

OpenStudy (aaronq):

sorry i screwed up the brackets, it should read: q(water)=-q(metal)

OpenStudy (anonymous):

so that is the specific heat?

OpenStudy (aaronq):

nope, you gotta solve the equation using the data of the metal.

OpenStudy (aaronq):

specific heat capacity is basically a measure of how much a certain substance retains heat.

OpenStudy (anonymous):

so is it something like: 685 ------- 26 x 6.3

OpenStudy (aaronq):

you have the concept right, but the change in temperature of the metal is not the same as the change in temp of the water.

OpenStudy (aaronq):

and you're using the mass of the metal not the mass of the water

OpenStudy (anonymous):

ohh so 27.776 --------- 26 x 6.3

OpenStudy (aaronq):

not quite, write in the units, it's easier to see where you went wrong. q=m*C*dT -> solve for C -> C=q/(m*dT) btw, dT=Tf-Ti

OpenStudy (anonymous):

but im not using C in specific heat am i?

OpenStudy (aaronq):

C of the metal is what the question is asking you to find

OpenStudy (anonymous):

lol oh yeeeah ok so hang on lemme try to figure this out

OpenStudy (aaronq):

hah okay

OpenStudy (anonymous):

685 ----------- 27.776 x 6.3

OpenStudy (aaronq):

it's right, except the change in temperature of the metal, it went from 100.5 to 31.6, clearly more than 6.3

OpenStudy (anonymous):

OMG I I ACTUALLY GET IT!!!

OpenStudy (anonymous):

so after i do this we're done?

OpenStudy (aaronq):

yes, but make sure you include the units at the end

OpenStudy (anonymous):

what would be the unit for specific heat??

OpenStudy (aaronq):

when you include the units in the equation you'll see because they won't cancel out

OpenStudy (anonymous):

oh ok thanks!!!

OpenStudy (aaronq):

no problem dude !